# CAP[Uncaptioned image] : Evaluation of Persuasive and Creative Image Generation

### v1

@@ Abstract
We address the task of advertisement image generation and introduce three evaluation metrics to assess Creativity, prompt Alignment, and Persuasiveness (CAP) in generated advertisement images. Despite recent advancements in Text-to-Image (T2I) generation and their performance in generating high-quality images for explicit descriptions, evaluating these models remains challenging. Existing evaluation methods focus largely on assessing alignment with explicit, detailed descriptions, but evaluating alignment with visually implicit prompts remains an open problem. Additionally, creativity and persuasiveness are essential qualities that enhance the effectiveness of advertisement images, yet are seldom measured. To address this, we propose three novel metrics for evaluating the creativity, alignment, and persuasiveness of generated images. Our findings reveal that current T2I models struggle with creativity, persuasiveness, and alignment when the input text is implicit messages. We further introduce a simple yet effective approach to enhance T2I models’ capabilities in producing images that are better aligned, more creative, and more persuasive. Code is available at https://github.com/aysanaghazadeh/CAP.
@@ 1 Introduction
Advertisements (ads) impact consumer decisions and appear across various media channels, from billboards to social platforms. Effective ads require more than a set of relevant objects; they must convey messages with creativity and persuasiveness to gain attention and persuade the audience to take an action. For example in Fig. 1, images (a), (b), (c) feature Gatorade. However, image (b) only shows the a bottle and two cans, while image (c) includes a creative slogan and an athlete running while holding it, and is more likely to persuade the audience to buy Gatorade.
@@
Figure 1: Human-created ads (a, d) convey their message through creative and persuasive visual storytelling, blending the implicit message seamlessly into the visuals. The T2I baseline (b, e), on the other hand, focuses on generating high-quality images by directly incorporating the action-reason prompt, often missing the underlying intent. This highlights the need for better metrics to measure persuasiveness, creativity, and abstract text-image alignment. Additionally, Ours (c, f) demonstrates how improving visual storytelling can enhance ad generation by aligning visuals with the intended message more effectively.
@@
Recent advances in text-to-image (T2I) models, such as Stable Diffusion [37, 33] and DALLE3 [2], excel at generating high-quality photorealistic images from explicit and detailed prompts that describe objects, their attributes and relations. However, a user of a T2I method for ad generation might desire to generate an ad that conveys a certain idea and thus use more visually implicit or abstract prompts (i.e. ones that do not precisely describe how the image should look in terms of objects), similar to the action-reason statements explored in [18, 27, 3] that use a reason (e.g., “it would help me win”) to persuade the viewer to take an action (e.g., “drink Gatorade”).
@@
Further, existing T2I generation metrics, like FID [15] and CLIPscore [14], assess general image fidelity and explicit text-image alignment but do not capture qualities that would be important for evaluating ads, such as creativity and persuasiveness. Our work addresses this gap by proposing the first benchmark on creative and persuasive ad generation, using three novel metrics: Creativity, text-image Alignment, and Persuasiveness (CAP).
We define Creativity as the divergence of the generated image from typical product images while maintaining alignment with the advertisement message. We propose a method to evaluate the Alignment of generated images with abstract messages, called Contextual Image-Text alignment Evaluator (CITE), by training an image description method to focus on the reason provided by generated ads, and comparing to the intended reason. Finally, we define Persuasiveness as being convincing in driving the action in the advertisement message through the intended reason, while also incorporating well-known strategies [36, 43, 40]. Our proposed alignment, creativity, and persuasiveness metrics achieve higher agreement with human judgment than baseline metrics, by 0.40, 0.51, and 0.59 (out of 1).
Our findings reveal that public T2I models struggle to generate creative and persuasive images aligned with the action-reason statements (AR). To explore this gap, we compare the generated images with real advertisement images, highlighting significant differences in creativity, persuasiveness, and alignment.
@@
To address the identified T2I shortcoming, we propose a simple yet effective method for the generation of advertisement images using LLMs to create advertisement image descriptions ($\mathcal{D_{LLM}}$). These descriptions are then used as prompts for the T2I models to generate advertisement images. Fig. 1 (c), produced by this method, successfully captures both the explicit and implicit aspects of the advertising message by including a slogan and showing an athlete running and holding the bottle of Gatorade. This simple approach improves the alignment, creativity, and persuasiveness by (6%, 10%, 12%) and (34%, 40%, 25%) in commercial ads and public service announcements, respectively.
@@
  • We introduce three novel evaluation metrics for advertisement image generation task: Creativity, Alignment, and Persuasiveness (CAP).
  • We show that current state-of-the-art public T2I models perform poorly in terms of creativity, persuasiveness, and alignment when generating advertisement images from abstract text, especially for public service ads.
@@ 2 Related Work
Text-image alignment. With the advancement of T2I models [11, 33, 5] in generating high-quality images, the evaluation of text-image alignment for explicit text descriptions has become a widely explored area. [17, 31] have introduced benchmarks for alignment evaluation. Unlike the advertisement generation task, the text prompts in both benchmarks are visually explicit descriptions. Some approaches compute image-image similarity between the ground-truth images (real images in a dataset) and the generated images, using CLIP [34] or FID [15]. These methods reward high similarity of the generated images to the ground-truth images. However, this evaluation overlooks the creativity of the images. Another group of metrics computes text-image alignment between prompt and generated images, using CLIP-score [14] and BLIP-score [23]. However, these methods fail even in evaluation of images generated for complex explicit prompts [25]. Some methods use LLMs and MLLMs to compute image-text alignment [25, 46, 42] by fine-tuning an LLM or MLLM to answer questions, e.g. “Does the image show [T2I Prompt]? Others [8, 51, 16] first generate questions given the T2I prompt and use these as prompts for VLMs to score the images. These methods have limited usability in our setting where the prompt is a message that does not explicitly mention the visual content, and two totally visually different images can convey the same message. Other methods [50, 48, 19, 20] rely on human feedback. [50] utilizes fine-tuned BLIP [23] as the backbone followed by a Multi-layer Perceptron to predict the score. [48, 20, 19] fine-tune CLIP [34] on the human feedback to predict the score given the prompt and image. Although these models are fine-tuned with (expensive) human feedback, they rely on VLMs, which have limited capacity to capture semantics in persuasive images [27]. Finally, [24, 12] focus on the image generation of concepts like peace rather than conveying a persuasive message and [47] proposes a method to generate images based on abstract description which still includes the visual elements. We propose a method to evaluate generation based on visually implicit text messages (few objects specified).
@@
Persuasion. Generation and evaluation of persuasive text-only content has been explored in the field of NLP. With the development of LLMs, many works [10, 45, 30] compare the persuasion in LLM-generated content with contents written by humans. [35] evaluates the performance of GPT-4 in understanding if the text content is convincing. [41] introduce benchmarks on persuasiveness of texts. [53, 29] propose methods to convert a text into a persuasive one or generate a persuasive text for specific prompt. [26] introduce a dataset on multimodal argumentative content. [18] introduce a dataset of advertisements which aim to be persuasive. Many non-computational works [36, 4, 32, 38] introduce and analyze different persuasion strategies in advertisement and their influence on effectiveness. [13, 22] introduce methods and benchmark for image persuasion strategy detection. However, to the best of our knowledge we are the first to introduce a computational method to score the level of persuasiveness of advertisement images, i.e. not how it is designed and what strategies it uses to try to be persuasive, but whether it is persuasive and creative. There are very few works focusing on generation of persuasive images; for example, [44] is limited to faces in ads.
Creativity. Creativity is essential in designing rhetoric content to make ads more effective. [39, 1] analyze the influence of creativity in ads. While [40] proposes a set of questions for analyzing creativity, it does not provide a computational metric; our work addresses this gap.
@@ 3.1 Background: Task Definition
We define advertisement generation as the creation of creative and persuasive advertisement images $(I_{gen})$ based on implicit input messages. The input messages ($AR_{m}$) and corresponding ground-truth images ($I_{real}$) are sourced from the PittAd dataset [18]. Each image in the PittAd dataset includes three to five action-reason messages ($AR_{m})$, which are the interpretation of the images structured as “I should {action}$(A_{m})$, because {reason} $(R_{m})$”.
Figure 2: Overview of CITE. Orange show training, while blue is inference. $AR_{w}$ and $AR_{l}$ are used in training as the preferred and dis-preferred statements. $AR_{m}$ is the prompt of the T2I model.
@@ 3.2 Alignment of Image with Abstract Text
In the advertisement image generation task, the input text often implies rather than explicitly describing the objects and their attributes. In these cases, generated images may appear visually relevant but fail to represent the deeper semantics or messages of the ad. For example, in Fig. 1 (b), the image includes a bottle of Gatorade, aligning visually with the input text, yet it fails to capture the emphasis on the reason “it would help me win” . While existing alignment evaluation methods [14, 50, 25, 51, 19] can detect visual mismatches, they struggle to capture higher-level semantic misalignment between $AR_{m}$ and the image. To address this, we introduce Contextual Image-Text alignment Evaluator (CITE), a metric to capture both visual and semantical mismatch. CITE evaluates the alignment of images with the implicit text by first utilizing a Multimodal Large Language Model (MLLM) to generate a detailed description of the images ($D_{I}$ in the Fig. 2). Next, we input this description into a fine-tuned LLAMA3-instruct-8B [9] to generate an action-reason statement for the input image ($AR_{gen}$ in the Fig. 2). We separately compute the similarity of the action and reason components of the $AR_{m}$ with the corresponding component in the $AR_{gen}$ using a semantical similarity score [6] and then return their weighted average as the overall alignment score, as follows:
$CITE(I_{gen},AR_{m})=\frac{CITE(I_{gen},A_{m})+\alpha\cdot CITE(I_{gen},R_{m})}{1+\alpha}$ (1)
@@
where $\alpha$ is the weight for reason when computing the similarity between the statements, which is set on 4 after experimenting with different values. $A_{m}$ and $R_{m}$ is action and reason components of the $AR_{m}$ respectively.
@@
While LLAMA3-instruct generates acceptable action-reason statements with an accurate prompt, it sometimes fails to generate statements with the correct reason given an image. For example, the LLAMA3-instruct output for Fig. 2 (b) is I should drive with confidence because Subaru promises a safe and reliable experience with its vehicles. While this is an acceptable output, driving with confidence (possibly over-confidently) is not what the image is advertising (being able to safely stop with confidence).
@@
To generate more accurate interpretations for the described images, and thus enable accurate alignment predictions, we fine-tune LLAMA3-instruct with Contrastive Preference Optimization (CPO) [49]. In CPO, each data point includes a prompt, preferred answer, and dis-preferred answer. The objective is to optimize the LLM to generate responses closer to the preferred answers while staying more distant from the dis-preferred answers. We fine-tune LLAMA3-instruct model by generating descriptions $(D_{I})$ for 250 images from [18] using InternVL-V2-26B [7]. These descriptions were then used as prompts during the training process. Then, we used each of the ground-truth action-reason statements from [18] as preferred action-reason statements $(AR_{w})$. To make the dis-preferred statements visually similar to actual statements and semantically different, we used each of the hard negatives from [27] as dis-preferred action-reason statements $(AR_{l})$. Hard negative statements share the visual elements with $AR_{m}$; however, the statements are semantically different. This generated a dataset in size of 3500 data-points. Finally, we fine-tuned LLAMA3-instruct using CPO on $(D_{I},AR_{w},AR_{l})$. The overview of the CITE process is shown in Fig. 2. After generation of $AR_{gen}$ by fine-tuned LLAMA3-instruct, we compute the similarity [6] between the generated action-reason statement $(AR_{gen})$ and the input message $(AR_{m})$.
@@ 3.3 Creativity
Prior work [40] defines creativity as uniqueness of the image. While uniqueness can make the image more creative and effective, it might result in irrelevant images to the text ($AR_{m}$). Based on the creativity definition and alignment criteria, we define creativity in the context of advertisement image generation as (1) being distant from an image which simply shows the objects in the message, and (2) conveying the message. Thus, the probability of the image being unique increases when it includes additional visual elements not explicitly mentioned in the prompt.
We introduce a creativity score as the text-image alignment score between the prompt and image (expected to be high) divided by the cosine similarity between the CLIP features of the image and the objects in the text, $sim(I_{gen},obj$), which is expected to be low. We formulate the creativity score ($C_{obj}$) as follows:
$C_{obj}=\frac{1}{n}\times\Sigma_{obj\in objects}\frac{CITE(AR_{m},I_{gen})}{sim(I_{gen},obj)}$ (2)
where $n$ is the number of objects mentioned in the text. To find the list of objects mentioned in the prompt, we utilize an LLM to find the object in text. As more elements become visible in the image, the similarity between the image and the objects in the text decreases. To ensure these additional elements do not detract from relevance and the image continues to convey the intended message, we incorporate CITE alignment in the score.
Figure 3: Process of computing PA score

### v2

@@ Abstract
We address the task of advertisement image generation and introduce three evaluation metrics to assess Creativity, prompt Alignment, and Persuasiveness (CAP) in generated advertisement images. Despite recent advancements in Text-to-Image (T2I) methods and their performance in generating high-quality images for explicit descriptions, evaluating these models remains challenging. Existing evaluation methods focus largely on assessing alignment with explicit, detailed descriptions, but evaluating alignment with visually implicit prompts remains an open problem. Additionally, creativity and persuasiveness are essential qualities that enhance the effectiveness of advertisement images, yet are seldom measured. To address this, we propose three novel metrics for evaluating the creativity, alignment, and persuasiveness of generated images. We show that current T2I models struggle with creativity, persuasiveness, and alignment when the input text is implicit messages. We further introduce a simple yet effective approach to enhance T2I models’ capabilities in producing images that are better aligned, more creative, and more persuasive. Code is available at https://aysanaghazadeh.github.io/CAP/
@@ 1 Introduction
Advertisements (ads) impact consumer decisions and appear across various media, from billboards to social platforms. Persuasive, effective ads satisfy some or all of the following three criteria: (1) convey their message clearly, (2) convey their message in a creative and memorable way, and (3) impact their intended audience (e.g. convince the audience to buy a product). Thus ads must do more than simply show a set of relevant objects. For example, in Fig. 1, images (a), (b), (c) feature Gatorade. However, (b) only shows a bottle and two cans, while (c) includes a creative slogan and an athlete running while holding the bottle, and is more likely to persuade the audience to buy Gatorade. Similarly, on the topic of child abuse, (e) shows a child, while (f) also shows a relevant ominous shadow.
@@
Figure 1: Human-created ads (a, d) convey their message through creative and persuasive visual storytelling, blending the implicit message seamlessly into the visuals. The T2I baseline (b, e), on the other hand, depicts relevant entities but without the underlying intent. This highlights the need for better metrics for persuasiveness, creativity, and abstract text-image alignment. Ours (c, f) demonstrates how improving visual storytelling can enhance ad generation by aligning visuals with the intended message effectively. Texts on the right form the prompt to T2I models.
@@
Recent advances in text-to-image (T2I) models, such as Stable Diffusion [38, 35] and DALLE3 [4], excel at generating high-quality photorealistic images from explicit and detailed prompts that describe objects, their attributes and relations. However, a user of a T2I method for ad generation might simply know what message they want the ad to convey, and may not have thought about the exact objects the image should depict. Thus, they may use visually implicit or abstract prompts (i.e. ones that do not precisely describe how the image should look in terms of objects).
@@
Further, existing T2I generation metrics, like FID [18] and CLIPScore [17], assess image fidelity and explicit text-image alignment but do not capture qualities that are important for evaluating ads, such as creativity and persuasiveness. Our work addresses this gap by proposing the first benchmark on creative and persuasive ad generation, using three novel metrics: Creativity, text-image Alignment, and Persuasiveness (CAP). These metrics correspond to the three criteria for effective ads that we posit above.
To measure how well the ad satisfies the first criterion (conveying the message clearly), we propose a method to evaluate the Alignment of Image and Message (AIM), by (1) training an LLM to interpret the argumentation (message) in the ad image, and (2) comparing the message inferred based on the generated image, to the intended message.
To measure satisfaction of the second criterion, conveying the message in a creative and memorable way, we define Creativity as the divergence of the generated image from typical product images while maintaining alignment with the ad message. Measuring satisfaction of the third criterion (impacting the intended audience) is more complex as we must consider both the ad content and the audience, through the lens of argumentation. We define Persuasiveness as being convincing in driving the message of the advertisement message to the intended audience while also incorporating well-known argumentation strategies [37, 45, 42].
To ground our work, we leverage [21], which gathers annotations similar to the implicit prompts that users might employ to generate ad images. These annotations are in the form of action-reason statements; they contain the action the ad suggests the viewer should take (e.g., “drink Gatorade”) and the reason the ad provides for taking this action (e.g., “it would help me win”). We use this dataset as a source of real ads and ad action-reason messages based on which we generate additional ads. On this data, our proposed alignment, creativity, and persuasiveness metrics achieve higher agreement with human judgment than baseline metrics, by 0.37, 0.50, and 0.53 (out of 1).
@@
Further, we find that public T2I models struggle to generate creative and persuasive images aligned with the action-reason statements. To address this T2I shortcoming, we propose a simple yet effective method to create expanded advertisement messages ($\mathcal{E_{LLM}}$) which use LLMs’ reasoning ability to explicitly describe what visual content an ad for the corresponding messages should contain. These expanded messages are then used as prompts for the T2I models to generate ads. Fig. 1 (c), produced by this method, successfully captures both the explicit and implicit aspects of the advertising message by including a slogan and showing an athlete running and holding the bottle of Gatorade. This simple approach improves alignment, creativity, and persuasiveness by (6%, 7%, 10%) and (39%, 38%, 43%) in commercial ads and public service ads, respectively. These numbers correspond to one particular configuration of MLLM, LLM and T2I methods involved, and we conduct extensive experiments with many configurations.
@@
  • We introduce three novel metrics for advertisement image generation: Creativity, Alignment, and Persuasion.
  • We show that current state-of-the-art public T2I models perform poorly in terms of creativity, persuasion, and alignment when generating advertisement images from abstract text, especially for public service ads.
@@ 2 Related Work
Text-image alignment. With the advancement of T2I models [13, 35, 7] in generating high-quality images, the evaluation of text-image alignment for explicit text descriptions has been widely explored. Some approaches compute image-image similarity between the ground-truth images (real images in a dataset) and the generated images, using FID [18] and IS [40]. These methods reward high similarity of the generated images to the ground-truth images. However, this evaluation over-focuses on low-level perceptual quality while overlooking the creativity of the images and alignment with intended messages. [^1] We experimented with FID and IS, but obtained counterintuitive results: better, more recent models like FLUX underperformed older models like SDXL, demonstrating these metrics are not appropriate for ad image generation evaluation. Another group of metrics computes text-image alignment between prompt and generated images, using CLIPScore [17] and BLIPScore [26]. However, due to the limitation of these methods in understanding compositional and complex prompts, these methods fail even in evaluation of images generated for complex explicit prompts [28]. Some methods use LLMs and MLLMs to compute image-text alignment [28, 49, 44] by fine-tuning an LLM or MLLM to answer questions, e.g. “Does the image show [T2I Prompt]?” Others [10, 54, 20] first generate questions given the T2I prompt and use these as prompts for VLMs to score the images. These methods have limited usability in our setting where the prompt is a message that does not explicitly mention the visual content, and two totally visually different images can convey the same message. Other methods [53, 51, 22, 23] rely on human feedback. Although these models are fine-tuned with (expensive) human feedback, they rely on VLMs, which have limited capacity to capture semantics in persuasive images [30]. Finally, [27, 14] focus on the image generation of concepts like peace rather than conveying a persuasive message and [50] proposes a method to generate images based on abstract description which still includes the visual elements. We propose a method to evaluate generation based on visually implicit text messages (few objects specified).
Non-computational analysis of persuasion and creativity. Theories of persuasion date to Aristotle’s pathos, ethos and logos [6]. Strategies specifically for decoding ads were proposed in [48]. Others [37, 6, 34, 39] analyze persuasion strategies and their influence on effectiveness and impact on the audience [45]. Finally, [41, 1, 42] analyze the influence of creativity in ads but do not propose computational metrics; our work addresses this gap.
@@
Computational analysis of persuasion. Generation and evaluation of persuasive text-only content has been explored in the field of NLP. With the development of LLMs, many works [12, 47, 33] compare the persuasion in LLM-generated content with contents written by humans. [36] evaluates the performance of GPT-4 in understanding if the text content is convincing. [43] introduce benchmarks on persuasiveness of texts. [56, 32] propose methods to convert a text into a persuasive one or generate a persuasive text for specific prompt. [29] introduce a dataset on multimodal argumentative content. [21] introduce a dataset of advertisements which aim to be persuasive. [16, 25] introduce methods and benchmarks for detecting image persuasion strategies like emotion, and reasoning in advertisements. However, to the best of our knowledge we are the first to introduce a computational method to score the level of persuasiveness of advertisement images, i.e. not classify the types of strategies it uses to try to be persuasive, but whether it is persuasive and creative. There are very few works focusing on generation of persuasive images; for example, [46] is limited to faces in ads.
@@ 3.1 Background: Task Definition
[…]
Figure 2: Overview of AIM. Orange denotes training, while blue is inference. $AR_{w}$ and $AR_{l}$ are used in training as the preferred and dis-preferred statements. $AR_{m}$ is the prompt for the T2I model.
@@ 3.2 Alignment of Image with Abstract Text
In the advertisement image generation task, the input text often implies rather than explicitly describing the objects and their attributes. In these cases, generated images may appear visually relevant but fail to represent the deeper semantics or messages of the ad. For example, in Fig. 1 (b), the image includes a bottle of Gatorade, aligning visually with the input text, yet it fails to capture the emphasis on the reason “it would help me win”. Existing alignment evaluation methods [17, 53, 28, 54, 22] fail to detect mismatches between messages ($AR_{m}$) used as prompts, and generated images, due to the abstractness (visual implicitness) of the messages. To address this, we introduce Alignment of Image and Message (AIM). We utilize two steps, shown in Fig. 2. First, we train (fine-tune) an LLM (e.g. LLAMA3-Instruct-8B) to match a description of the ad image to the correct action-reason statement, ensuring the LLM pays fine-grained attention to the semantics of the message. Second, we use this trained LLM to produce an action-reason statement for a generated image ($AR_{gen}$), and compute the alignment (similarity) between this produced action-reason statement and the input action-reason statement in the prompt ($AR_{m}$), as the AIM score for the generated image.
In more detail, we utilize a Multimodal Large Language Model (e.g. InternVL-V2-26B [9]) to generate the image description $D_{I}$. Next, we input this description into a fine-tuned LLM (e.g. LLAMA3-Instruct-8B [11]) to generate an action-reason statement for the input image ($AR_{gen}$). We separately compute the similarity of the action and reason components of the $AR_{m}$ (i.e. $A_{m}$ and $R_{m}$) with the corresponding component in the $AR_{gen}$ (in order to avoid the more visually explicit action from dominating the score). We use a semantical similarity score [8] and then return their weighted average as the overall alignment score:
$\small AIM(I_{gen},AR_{m})=\frac{Sim(A_{gen},A_{m})+\alpha\ Sim(R_{gen},R_{m})}{1+\alpha}$ (1)
@@
where $\alpha$ is the weight for reason when computing the similarity between the statements, which is set [^2] Agreement between human annotators and AIM scores computed with different $\alpha$, evaluated on a subset of images, are: $\alpha=1:0.51$, $\alpha=2,3:0.55$, $\alpha=4,5:0.68$. We include an example in the supplementary file. to 4.
@@
We opt to fine-tune the LLM because it sometimes fails to generate statements with the correct message given an image. For example, the LLAMA3-Instruct-8B output for Fig. 2 (b) is I should drive with confidence because Subaru promises a safe and reliable experience with its vehicles. While this is an acceptable output, driving with confidence (possibly over-confidently) is not what the image is advertising (being able to safely stop with confidence).
@@
To generate more accurate interpretations for the described images, and thus enable accurate alignment predictions, we fine-tune an LLM with Contrastive Preference Optimization (CPO) [52]. In CPO, each data point includes a prompt, preferred answer, and dis-preferred answer. The objective is to optimize the LLM to generate responses closer to the preferred answers while staying more distant from the dis-preferred answers. The prompts are image descriptions $(D_{I})$ for 250 images from [21] (generated using InternVL-V2-26B [9]). Each ground-truth action-reason statement from [21] serves as a preferred action-reason statement $(AR_{w})$. For dis-preferred action-reason statements $(AR_{l})$, we used the hard negative statements from [30]. These hard negatives share the visual elements with $AR_{m}$; however, the statements are semantically different. For example, in Fig. 2, “I should buy this car because it is a race car on the inside” is converted to “I should buy this car because it lacks the performance of the race car on the inside”. While the visual elements—car and race car—are the same in both sentences, the semantics of the sentence are completely changed. This process generated a dataset in size of 3500 data points. Finally, we fine-tuned the LLM using CPO on $(D_{I},AR_{w},AR_{l})$, as shown in Fig. 2. Recall that after generation of $AR_{gen}$ by the fine-tuned LLM, we compute the similarity [8] between the generated action-reason statement $(AR_{gen})$ and the input message $(AR_{m})$.
@@ 3.3 Creativity
We define creativity in the context of advertisement image generation as (1) uniqueness, i.e. being distant from an image which simply shows the objects in the message, while maintaining (2) relevance, i.e. conveying the intended action-reason message. We posit that the probability of the image being unique should increase when it includes additional visual elements not explicitly mentioned in the prompt. Thus, we introduce a creativity score with two components: (1) cosine similarity between the CLIP features of the image and the objects in the text, $Sim(I_{gen},obj$) (expected to be low for unique images); and (2) text-image alignment between the prompt/message and image (expected to be high for relevant images). We formulate the creativity score ($C_{obj}$) as follows:
$\small C_{obj}=\frac{AIM(I_{gen},AR_{m})}{\frac{1}{n}\times\Sigma_{obj\in objects}~Sim(I_{gen},obj)+0.01}$ (2)
where $n$ is the number of objects mentioned in the text, and 0.01 is a smoothing value to avoid diving by 0. The smoothing value is chosen to be small compared to average CLIP scores (which is about 0.30 in our experiments) to limit the effect of soothing value. We use an LLM to extract a list of objects mentioned in the action-reason text.
Figure 3: Process of computing $P_{comp+AIM}$ persuasiveness score